two towns are 1300 miles apart. A group of hikers start from each town and walk the trail toawrd each other. They meet after a total hiking time of 200 hours. If one group travels 2 1/2 mile per hour faster than the other group. find the rate of each group.
Answer by checkley79(3341)

D=RT
1300=(X+2.5X)200
1300=3.5X*200
1300=700X
X=1300/700
X=1.857 MPH. FOR THE SLOWER GROUP.
2.5*4.64 MPH. FOR THE FASTER GROUP.
PROOF:
1300=(1.875+4.64)200
1300=6,515*200
1300=1300=1300